[figure: 9.7 cm]

INTERFACE, MISFIT AND BULK MAGNETOELASTIC COUPLINGS IN RARE EARTH (Dy, Ho, Er)/Y, Lu, Tm SUPERLATTICES

A. del Moral^{∗}, J.I. Arnaudas^{∗}, M. Ciria^{∗}, C. de la Fuente^{∗}, L. Benito^{∗},

M.R. Wells^{#} and R.C.C. Ward^{#}

(∗) Dpto. Magnetismo de Solidos, Dpto. Fisica de Materia Condensada & ICMA,

Universidad de Zaragoza-CSIC, 50009-Spain (#) Dpt. of Physics, Clarendon Laboratory, Oxford University, OX1 3PU, UK

We review our researches on the magnetoelastic (MEL) coupling (MEL-C) in a wide range of MBE grown Rare Earth (RE = Ho, Dy, Er) superlattices (SL), spaced by Y, Lu and Tm (SP), of formula RE_{n}/SP_{m}. The epitaxy is [11\overline{2}0]Al_{2}O_{3}∥[110]Nb∥[0001]RE/[110]Nb, and RE blocks are hcp pseudomorphic. We performed MEL stress (MEL-S) measurements up to 12 T fields, B, between 10-300K. MEL-S σ̃(a, a) and σ̃(a, b), were measured on basal plane (b.p.), with B ∥ easy a-direction (Dy, Er), b(Ho), using a capacitive cantilever technique. We determined the irreducible M^{γ} MEL-S, breaking the b.p. hexagonal symmetry. The series are: Ho/Lu (n=8–85; m=15), Ho/Y (10; 10–40, 6; 6), Dy/Y (5-25; 15); Er/Lu (10-30; 10); Ho/Tm (8-60; 16). MEL-S for thick (10^{4} Å) RE films gave close values to the bulk. The MEL-S in RE’s SL are rather different as in bulk. We found three contributions to M^{γ} : RE volume (M{γ \atop v0}), interface (IF) (M{γ \atop s}) and RE/SP a(or b)–lattice constant misfit (MF) strain ϵ{γ \atop mf}, i.e. we observed second-order MEL effects. They compose as: M^{γ} = M{γ \atop v0}+2M{γ \atop s} /t_{RE} +D{γ \atop v} ϵ{γ \atop mf}, for elastic regime (ϵ{γ \atop mf} = e(t_{SP} /(α t_{RE}+t_{SP}))), and bulk mismatch e= (a{b \atop SP} - a{b \atop RE})/a{b \atop SP} = -1.7%; in Ho/Lu SL, Ho is under compression. Edge dislocation effects could also be present. The main results, e.g. in the Ho_{n}/Lu_{15} SL, are:

i) the IF and MF MEL-S are strong, -6 and 5.5 times M{γ \atop v0} (n=8), with MEL-S (in GPa and at 0K) M{γ \atop v0} =+0.275, M{γ \atop s} /(c/2) =-7.0 and D{γ \atop v} =-116; ii) MEL-S are of single–ion crystal-field (CEF) origin, as concluded from: a) scalings with reduced magnetisation, m(T), are as Î_{5/2}(L^{−1}(m)) for M{γ \atop v0} and D{γ \atop v}, and as m^{4} for M{γ \atop s}, i.e. RE-IF behaves as 2d − spin; b) the MEL-S signs agree with Stevens parameter α_{J}. We also studied Ho_{10}/Y_{n} SL to single out the MF MEL-S; M^{γ}/(c/2) changes sign at n ≅ 14, because of the opposite signs of M{γ \atop v0}=0.27 GPa and M{γ \atop s} /(c/2) = 2.16 GPa relative to D{γ \atop v} = -56 GPa. In Ho/Y SL e=+2.3%, so Ho is under tension. In Ho_{n}/Tm_{16} SL, with competing anisotropies ions, Ho^{3+} and Tm^{3+},we found a sign inversion of M^{γ} relative to Ho_{n}/Lu_{15} SL. A developed model shows that the strong and of opposite sign relative to M{γ \atop v0}, M{γ \atop s} comes out from localised conduction band states at IF, screening the IF distorted CEF differently as in the volume. For revealing the α − mode MEL-S (of volume M{α \atop 12} and of c/a distortion, M{α \atop 22}) we studied the thermal variation (4.2K-300K) of a and c parameters for Er_{n}/Lu_{10} SL, using a triple-axis X-ray spectrometer. Noticeably, besides the exchange α − mode, we found strong IF and MF CEF contributions. Finally, accurate magnetic phase diagrams in (H_{C} ,T) plane were traced from our MEL-S and magnetization measurements.

[figure: 9.7 cm]
[figure: 13.4 cm]

Subject category :

2. Magnetic Films, Surfaces and Multilayers

Presentation mode : PLENARY LECTURE

Corresponding author : A. del Moral

Address for correspondence : Dpto. Magnetismo de Slidos, Dpto. Fsica de Materia Condensada & ICMA, Universidad de Zaragoza-CSIC, 50009-Spain

Email address : delmoral@posta.unizar.es